1a) Consider an economy where output is produced using just labor as input. Output is sold at $10 per unit and the marginal product of labor for each worker hired is expressed as MPL=10-2L. If each worker is paid $40, what would be the total profit of the profit-maximizing entrepreneur?

a. $40

b. $60

c. $0

1b) Consider an economy where output is produced using just labor as input. Output is sold at $10 per unit and the marginal product of labor for each worker hired is expressed as MPL=10-2L. How many workers should the profit-maximizing entrepreneur hire?

a. 1

b. 3

c. 4

1c) Consider an economy where production occurs without capital, and output Y sells for $10 a unit. Each worker hired is paid a wage of $20, and each worker is capable of producing 4 units of output. How many workers should the profit-maximizing producer hire?

a. The 5th worker

b. He should hire as many as he can find

c. The 10th worker
Answer #1

MPL = 10-22 W- $40 At Profit maximizing Condition UMPL = W VMPL = (P)(MPL) = (10) I 10-21) UMPL = 100 - 200 100- Qol = Yo theTC = (W) (L) = (40)(3) - $126 Total profit = 1R-TC har = $(120-130) - Hence option (C) is correct MPL 10-22 The profit maximiBecause umpl> Wage at each level of output or for each unit of Worker. Therefore, the propt maximising from should hire as ma